# Neural Combinatorial Optimization Algorithms for Solving Vehicle Routing Problems: A Comprehensive Survey with Perspectives

### v2

@@ VI-C Inadequacies of Vehicle Routing Problem Variants
As shown in Section II-A, most NCO solvers specialize in TSP and CVRP, exhibiting inadequacies when faced with VRP incorporating other constraints [192, 193]. Consequently, there arises a necessity to formulate solvers explicitly designed for VRP variants with intricate constraints in realistic scenarios [194, 195]. For example, dispatching vehicle fleets within the premise of an airport can be considered a multiple-fleet VRP [6], encompassing various constraints such as capacity, time window, precedence, etc. In this subsection, we investigate methodologies for customizing NCO solvers to address the constraints posed by MOVRP and DVRP, two prominent VRP variants entailing constraints in real-world scenarios.
To tackle the problem arising from the multiple objectives in MOVRP, certain studies [105, 106, 107, 108] adopt the MOEA/D framework [196]. This framework is designed to decompose a multi-objective problem into multiple distinct single-objective problems, and develops models corresponding to each single-objective problem, thereby addressing them individually. Finally, the Pareto set is constructed based on solutions of all single-objective problems. Furthermore, Lin et al. [110] proposed an approach to approximate the Pareto set using a single preference-conditioned solver, instead of employing multiple solvers. This preference-conditioned solver takes preference vectors (each vector denotes a single-objective problem) as inputs and constructs all possible (sub-)optimal solutions on the Pareto front.
@@
Recently, several studies [197, 198, 199] integrated Multi-task Learning into the field of NCO, enabling solvers to adeptly tackle a range of VRP variants. For example, Liu et al. [197] proposed a joint loss function comprising multiple sub-loss functions, each corresponding to a class of VRP variants, and all carrying equal weightage. However, learning multiple tasks simultaneously presents challenges not encountered in single-task learning, primarily due to potential conflicts between different task requirements [200, 201]. In the future, we advocate for alleviating conflicts by adjusting the weightage of the loss function for each task. For example, Meta Learning offers a viable solution, which can discern relationships between tasks and dynamically adjust their weightage. Once these powerful and versatile solvers have been developed to handle multiple VRP variants without requiring special designs and retraining, we anticipate the industry will be keen to adopt them in a wide range of applications.
@@ VII Conclusion
This paper presents a systematic review to assess existing NCO solvers and identify open challenges for future research. This work achieves the following accomplishments while prior surveys did not, namely 1) all existing NCO solvers can be categorized accordingly w.r.t our proposed taxonomy; 2) the inadequacies of existing NCO solvers, including generalization, as well as the resolution of large-scale and multi-constrained VRPs, are comprehensively introduced; 3) the advantages of SL and UL learning paradigms are well presented; and 4) the difficulties in applying NCO solvers in real-world VRPs are discussed. At the same time, we propose numerous promising research directions to address these inadequacies. It must be acknowledged that due to the wide array of VRP variants, this survey solely delineates the inadequacies specific to the major variants. Nevertheless, this survey can alleviate the difficulties faced by fellow researchers dealing with non-traditional VRP variants. In addition, we provide a live web page repository for NCO solvers organized according to our proposed taxonomy, continuously updating with emerging NCO solvers. We anticipate that this survey and the live repository will fulfill their intended objective of offering an overview of the latest developments in NCO solvers for VRPs, thereby serving as a valuable resource for researchers and practitioners.

### v3

@@ VI-C Inadequacies of Vehicle Routing Problem Variants
As shown in Section II-A, most NCO solvers specialize in TSP and CVRP, exhibiting inadequacies when faced with VRP incorporating other constraints [211, 212, 213]. Consequently, there arises a necessity to formulate solvers explicitly designed for VRP variants with intricate constraints in realistic scenarios [214]. For example, dispatching vehicle fleets within the premise of an airport can be considered a multiple-fleet VRP [6], encompassing various constraints such as capacity, time window, precedence, etc. In this subsection, we investigate methodologies for customizing NCO solvers to address the constraints posed by MOVRP and DVRP, two prominent VRP variants entailing constraints in real-world scenarios.
To tackle the problem arising from the multiple objectives in MOVRP, certain studies [111, 112, 113, 114] adopted the MOEA/D framework [215]. This framework is designed to decompose a multi-objective problem into multiple distinct single-objective problems, and develops models corresponding to each single-objective problem, thereby addressing them individually. Finally, the Pareto set is constructed based on solutions of all single-objective problems. Furthermore, Lin et al. [116] proposed an approach to approximate the Pareto set using a single preference-conditioned solver, instead of employing multiple solvers. This preference-conditioned solver takes preference vectors (each vector denotes a single-objective problem) as inputs and constructs all possible (sub-)optimal solutions on the Pareto front. In addition, certain studies have demonstrated the value of incorporating the imaging modality to address MOVRP. Chen et al. [216] proposed a graph-image bimodal fusion framework that enhances the performance of NCO by integrating graph and image information.
@@
Recently, several studies [217, 218, 219, 70, 220, 221] integrated Multi-task Learning into the field of NCO, enabling solvers to adeptly tackle a range of VRP variants. For example, Liu et al. [217] incorporated the concept of compositional zero-shot learning to propose MTL, which formulates VRPs as different combinations of a set of constraints (e.g., capacity) and exploits a shared network to encode these constraints. To train the solver, MTL employs a joint loss function comprising multiple sub-loss functions of equal weightage, each corresponding to a distinct VRP variant. Zhou et al. [70] incorporated the concept of Mixture-of-Experts to further enhance the performance of MTL. However, learning multiple tasks simultaneously presents challenges not encountered in single-task learning, primarily due to potential conflicts between different task requirements [222, 223]. We advocate for alleviating conflicts by adjusting the weightage of all the sub-loss functions. For example, to better address this problem, Berto et al. [221] proposed mixed batch training and multi-variant reward normalization. The former ensures that each training batch comprises VRP instances with diverse constraints, while the latter normalizes the reward for each VRP variant. In addition, Meta Learning offers a viable solution, which can discern relationships between tasks and dynamically adjust the weightages of their corresponding sub-loss functions. It is worth noting that certain studies [224, 225] have also explored using a versatile solver to address different COPs, such as VRP and JSP, a goal that poses greater challenges than designing a unified solver for VRP variants.
Currently, these versatile solvers are limited to handling VRPs of sizes up to 100 and cannot effectively handle instances from different distributions. Future research should focus on enhancing their performance on large-scale VRP instances and across various distributions, while progressively incorporating additional VRP variants. Once these powerful and versatile solvers have been developed to handle large-scale VRPs across diverse instance distributions and problem constraints, without the need of special designs and retraining, we anticipate the industry will be keen to adopt them in a wide range of applications.
@@ VII Conclusion
This paper presents a systematic review to assess existing NCO solvers and achieves the following accomplishments while prior surveys did not, namely 1) all existing NCO solvers can be categorized accordingly w.r.t our proposed taxonomy; 2) the inadequacies of existing NCO solvers, including generalization, and the resolution of large-scale and multi-constrained VRPs, as well as on-going efforts to solve these inadequacies are comprehensively introduced; 3) the advantages of SL and UL learning paradigms are well presented; and 4) the difficulties in applying NCO solvers in real-world VRPs are discussed. Through a comprehensive investigation, we identify open inadequacies for future research. Notably, on-going efforts focus on only one or two of these inadequacies, with none attempting to address all of them. Furthermore, we propose numerous promising research directions to address these inadequacies. It must be acknowledged that due to the wide array of VRP variants, this survey solely delineates the inadequacies specific to the major variants. Nevertheless, this survey can alleviate the difficulties faced by fellow researchers dealing with non-traditional VRP variants and the other COPs. In addition, we provide a live web page repository for NCO solvers organized according to our proposed taxonomy, continuously updating with emerging NCO solvers. We anticipate that this survey and the live repository will achieve their intended objective of offering an overview of the latest developments in NCO solvers for VRPs, thereby serving as a valuable resource for researchers and practitioners.
At the same time, research on NCO solvers provides significant benefits to other scholars from related fields such as the broad machine learning and specific OR communities. For example, to reorder the Optical Character Recognition (OCR) text blocks into a logical reading sequence, Li et al. [229] encoded each OCR text block as a node that encapsulates both visual layout and text semantic features, and subsequently employed a Ptr-Net-based network [29] to perform the reordering. To generate a shorter advertisement video, Tang et al. [230] have also drawn on NCO solvers and proposed M-SAN, which first segments the original video and subsequently selects and assembles certain segments into a shorter video. Furthermore, investigating the geometric invariance properties of VRPs is also valuable for GNN-based point cloud analysis methods [231]. Advancements in NCO solvers will surely benefit these relevant machine learning studies. In the OR community, several studies (e.g., [46, 50, 166]) have successfully integrated NCO techniques into traditional OR algorithms, improving their efficiency. For example, the classic LKH algorithm [58] for solving VRP problems is computationally intensive because it requires a non-negligible number of iterations to obtain node penalties. Moreover, LKH relies on manually specified rules for constructing edge candidate sets, which could limit solution quality. To address this problem, Xin et al. [46] proposed NeuroLKH to simultaneously predict the edge candidate sets and node penalties. Specifically, NeuroLKH begins by encoding edge and node inputs into feature vectors and then employs two distinct decoders to predict edge scores and node penalties, respectively. We anticipate that future research will further exploit the complementary strengths of OR and NCO to address different optimization problems.